A particular fruit's weights are normally distributed, with a mean of 263 grams and a standard deviation of 27 grams. The heaviest 6% of fruits weigh more than how many grams? Give your answer to the nearest gram.
Answer #1

Solution:-

Given that,

mean = = 263

standard deviation = =27

Using standard normal table,

P(Z > z) = 6%

= 1 - P(Z < z) = 0.06

= P(Z < z) = 1 - 0.06

= P(Z < z ) = 0.94

= P(Z < 1.56 ) = 0.94

z =1.56

Using z-score formula,

x = z * +

x = 1.56 * 27+263

x = 305.12

x=305